### v1

@@ Appendix A Appendix
PHX
True
True
False
False
False
30270-4768
20
PHX
True
True
True
False
False
30270-4484
16
PHX
True
True
False
True
False
30270-4363
17
PHX
True
True
False
False
False
30270-3842
14
SFO
True
True
True
False
False
30270-3838
13
PHX
True
True
False
False
False
30270-1730
6
PHX
True
False
False
False
False
30270-1583
7
PHX
True
True
True
True
False
30270-1501
@@
PHX
True
True
True
False
False
NA
4
PHX
True
True
True
False
False
NA
1
PHX
True
True
True
True
False
As stated in the Methods section (section 2), the “NHTSA SGO In-Transport Exclude Low delta-V (DV)” category was derived by applying a impulse-momentum collision model for vehicle-to-vehicle crashes. Each of the 7 ADS crashes with a fixed or non-fixed object was examined individually to estimate a delta-V. All but 2 of the 7 vehicle-to-object crashes were excluded from the low delta-V category. Three (3) crashes resulted in underbody damage when the Waymo vehicle “bottomed out” entering a parking lot driveway (30270-5612, 30270-5611) or tire damage after driving through a pothole (30270-5114) which resulted in less than the 1.0 mph threshold. Two (2) crashes involved non-fixed objects with low mass (<100 kg): a cardboard box (30270-5081) and a swinging parking lot gate that was not fixed (30270-4363).
Of the two vehicle-to-object crashes included, SGO report 30270-6548 involved a Waymo ADS vehicle that was driving in a construction zone and “entered a lane undergoing construction …, encountered a section of roadway that had been removed, and the front driver’s side wheel dropped off the paved roadway.” After reviewing the on-board data from this crash event, the front wheel dropping likely caused a delta-V of 1 to 3 mph. SGO report 30270-6561 involved a Waymo ADS vehicle “exited a parking lot…, when the passenger’s side sensors and side mirror made contact with an automatic gate that was in the process of closing.” Based on review of the sensor data, the contact with the closing gate caused damage to a sensor housing and side view mirror of the vehicle, which are unlikely to result in a delta-V above 1 mph. Because the event in report 30270-6561 was police-reported, we decided to not exclude this event from the low delta-V category.

### v2

@@ 14 Rate Ratio and (lower, upper) confidence intervals
20 point = (Y / t) / (X / s) if X > 0 else float("inf")
21 lower = s / t * betaprime.ppf(alpha / 2, Y, X + 1) if Y > 0 else 0
22 upper = s / t * betaprime.ppf(1.0 - alpha / 2, Y + 1, X)
23
24 return point, (lower, upper)
Listing 2 shows test cases based on select results from Table 7. Listing 3 shows the output of the test code.
Listing 2: Test Cases to Compute Nelson (1970) for Example Results from Table 7
1import sys
2import scipy
3print(f"Python version: {sys.version_info}")
[…]
@@
6# (Label, Benchmark IPMM, Benchmark Million Miles, Waymo Count,
7# Waymo Million Miles) from Table 2 (benchmark) and Table 4 (Waymo)
8phx_benchmark_m_miles = 24865
9sf_benchmark_m_miles = 862
10phx_waymo_m_miles = 5.343
11sf_waymo_m_miles = 1.755
12test_cases = [
13 ("Police-reported PHX", 4.31, phx_benchmark_m_miles, 12, phx_waymo_m_miles),
14 ("Police-reported SF", 5.86, sf_benchmark_m_miles, 3, sf_waymo_m_miles),
15 ("Any-injury PHX", 1.81, phx_benchmark_m_miles, 2, phx_waymo_m_miles),
16 ("Any-injury SF", 5.82, sf_benchmark_m_miles, 1, sf_waymo_m_miles),
17]
18
19# (Rate Ratio, CI Lower, CI Upper) as reported in the Table 7
20expected = [
21 (0.52, 0.24, 0.97),
22 (0.29, 0.05, 0.95),
23 (0.21, 0.02, 0.84),
24 (0.10, 0.001, 0.62),
25]
26
27# results are round to 2 decimal places
28test_tol = 1e-2
29for test_input, test_expected in zip(test_cases, expected):
30 (label, benchmark_ipmm, benchmark_m_miles, waymo_count, waymo_m_miles) = (
31 test_input)
32 (expected_point, expected_ci_lower, expected_ci_upper) = test_expected
33 benchmark_count = benchmark_ipmm * benchmark_m_miles
34 rate_ratio, (ci_lower, ci_upper) = nelson_ci(
35 benchmark_count, benchmark_m_miles, waymo_count, waymo_m_miles)
36 print(f"Comparison: {label}, Rate Ratio: {rate_ratio:.2f}, "
37 f"CI lower: {ci_lower:.2f}, CI upper: {ci_upper:.2f}")
38 assert abs(rate_ratio - expected_point) < test_tol
39 assert abs(ci_lower - expected_ci_lower) < test_tol
40 assert abs(ci_upper - expected_ci_upper) < test_tol
Listing 3: Output of Test Cases
1Python version: sys.version_info(major=3, minor=11, micro=8, releaselevel=’final’, serial=0)
2scipy version:1.10.1
3
4Comparison: Police-reported PHX, Rate Ratio: 0.52, CI lower: 0.24, CI upper: 0.97
5Comparison: Police-reported SF, Rate Ratio: 0.29, CI lower: 0.05, CI upper: 0.95
6Comparison: Any-injury PHX, Rate Ratio: 0.21, CI lower: 0.02, CI upper: 0.84
7Comparison: Any-injury SF, Rate Ratio: 0.10, CI lower: 0.00, CI upper: 0.63
A.4 Considerations for Safety Impact Assessment for Serious Injury and Fatal Crashes
Conclusions regarding serious injury alone, or fatalities alone, require more data because they are subsets of the any-injury-reported outcome group which combines all injury and fatality results into one data category. ADS developers use multiple, complementary safety assurance methods to mitigate risk of high severity outcomes across architectural, behavioral, and operational layers (e.g., platform verification and validation, hazard analysis, see Webb et al. 2020, Favarò et al. 2023a). Traditionally, automotive safety systems have used prospective safety benefit analyses that use test track data and simulations as risk assessment tools and to determine potential high severity benefits. These types of prospective studies have also been performed for ADS. For example, one study that used reconstructions of all fatal crashes in Chandler, AZ found that a level 4 ADS was able to prevent 100% of crash when placed in the initiator role and prevent 82% and mitigate 10% when placed in the responder role (Scanlon et al., 2021).
Furthermore, ADS are designed to be able to follow applicable speed limits on the roads they drive on. Reducing speeding is a core pillar of the Vision Zero and Safe Systems approach to reducing serious injuries and fatalities on roadways. The Safe Systems approach has been adopted as a policy initiative by the U.S. Department of Transportation (United States Department of Transportation, 2022). The results of this study show that the ADS evaluated had a statistically significant lower any-injury-reported crashed vehicle rate. Serious injury and fatalities are a subset of this any-injury-reported benchmark, but no statement on these outcome levels can be made at this time based on this retrospective data. In a Safe Systems approach, it is not a requirement to prove efficacy of countermeasures before they are deployed in the field. Due to the statistical requirements, like the ones presented in Scanlon et al. (2023) or Kalra and Paddock (2016), assessing the safety impact of a system on serious injury or fatal crashes is impractical in many situations. The Safe Systems approach instead uses known energy-related factors that contribute to a large amount of serious injury and fatal crashes (e.g., excessive speed, or excessive energy transfer) and attempts to control these factors to stay below the tolerances to lead to these outcomes.
As ADS RO miles increase, the likelihood of observing a serious injury or fatality also increases. As discussed in the statistical power analysis of Scanlon et al. (2023), hundreds of millions to billions of miles of VMT would be necessary to detect a statistically significant difference in fatal crashed vehicle rates, depending on the magnitude of the difference in crashed vehicle rate of the ADS and benchmark. The national rate of crashed passenger vehicles involved in fatal crashes while traveling on surface streets is 22.3 incidents per billion miles (IPBM) (Scanlon et al., 2023). Consider a fictive driver who experiences 1 fatal crash after driving 10 million miles. At the time of this fictive fatal crash, the driver’s fatal crashed vehicle rate would be 100 IPBM with 95% confidence intervals of [3, 557] IPBM. Although the point estimate of 100 IPBM is higher than the benchmark, due to the low event count relative to the miles, this fictive driver’s crashed vehicle rate is not statistically different from the benchmark. Consider another fictive driver involved in 1 fatal crash in 100 million VMT, which is a rate of 10 IPBM. The 95% confidence intervals of this fictive driver is [0.3, 56] IPBM, which is still overlapping with the benchmark and not significantly different.